e^x times x^x is equals to e^2x. Is it true or not either true or false say the answer

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
e^x times x^x = e^x * x^x = (ex)^x
 
Hence it is not equal to e^(2x)
 
Hope it helps
 
Regards
Arun